OpenStudy (anonymous):

A line has slope –5/3. Through which two points could this line pass? (1 point)

OpenStudy (anonymous):

The possible answers are (12, 13), (17, 10) , (16, 15), (13, 10), (0, 7), (3, 10) & (11, 13), (8, 18) I cant seem to get it right.

OpenStudy (anonymous):

(12,13) (17,10) gets me -3/5 do i flip that to get -5/3?

OpenStudy (rob1525):

The question doesn't provide you with enough info you have to test the possible answers by using the point slope folmula y-y1=(-5/3)(x-x1).

OpenStudy (anonymous):

rob is right i dont see how you can get you answer just with that information, you at least need a y intercept

OpenStudy (anonymous):

here is the exact question : A line has slope –5/3. Through which two points could this line pass? (1 point) (12, 13), (17, 10) (16, 15), (13, 10) (0, 7), (3, 10) (11, 13), (8, 18)

OpenStudy (anonymous):

how can your points be positve if your rise is negative

OpenStudy (anonymous):

Im not sure :P

OpenStudy (rob1525):

so plug one point into the point slope form and the equation of the line then plug the x part from the other poin and see if the equation gives you the same y value.

OpenStudy (rob1525):

like this y-12=(-5/3)(x-13) now solve y=(-5/3)x(-5*-13/3)+12 which is y=(-5/3)x+(101/3) now plug in the x from the second point...17 and see if it equals 10.

OpenStudy (rob1525):

nope it equls 5. You see. do the same for the other 3 or until you get the right y.